# Covariate balancing estimation and model selection for difference-in-differences approach

### v1

@@ 3.2 Numerical experiments
The robustness of the proposed estimating method against model misspecification is examined by simulation. The settings for data generation follow
@@
Throughout this article, all random numbers are generated independently. The parameter $\beta^{*}$ represents the magnitude of the contribution of the covariate $x_{1}$ to the causal effect. The propensity score is estimated using only $x_{1}$, that is, the parameter $\alpha^{*}$ represents the magnitude of the contribution of $x_{2}$ to the true propensity score, which is not included in the model for estimating the propensity score. Note that when $\alpha^{*}\neq 0$, we are considering a model misspecification in which $x_{2}$ should be included in the model. Be that as it may, since (7) holds, the proposed method is theoretically guaranteed to be consistent. Under this setting, we compare the proposed method with a naive method estimating the propensity score by MLE, and confirm whether the proposed method is robust to the misspecification of the propensity score model by numerical experiments. The number of repetitions is 3000 in all numerical experiments of this article.
In Table 1, the true values of the ATT, the mean of the estimates and their empirical confidence intervals are given for several $\alpha^{*}$, $\beta^{*}$ and sample sizes $n$. With increasing $\alpha^{*}$, which represents the degree of the misspecification of the propensity score model, the bias of the naive method becomes larger, and the true value is no longer included in the confidence interval. Regarding the proposed method, it is found that the approximation accuracy is considerably high when the weighting matrix in the generalized method of moments is the identity matrix. On the other hand, when the weighting matrix is an optimal matrix, the accuracy is often comparable to that of the identity matrix, but the accuracy is similar to that of the naive method in some cases. Since there are settings in the simulation where the optimal matrix is close to degenerate, this may be attributed to the fact that the estimation of the $\theta$ is not stable. Regardless of whether or not (7) holds, if the propensity score model is estimated correctly, both the proposed method and the naive method can estimate ATT without bias.
Table 1: Comparison of the estimator of the ATT when the propensity score is given by the CBD method and the MLE: True is the true value of the ATT, CBD-id is the CBD method when the weighting matrix in the generalized method of moments is the identity matrix, and CBD-opt is the CBD method when it is the optimal matrix. Est is the average of the estimators obtained for each simulation, and Conf Int is its empirical 95% confidence interval.
@@
4 Model selection criteria for the semiparametric difference-in-difference method when propensity score is known
4.1 Risk function and its asymptotic evaluation in the semiparametric difference-in-difference method
As mentioned in the Introduction, there are no reasonable model selection criteria even for the basic setting as treated in Abadie (2005), we first derive model selection criteria when the propensity score is known, and then develop model selection criteria when using the proposed CBD method. Specifically, using the ideas of Baba et al. (2017) and Baba and Ninomiya (2021), we derive model selection criteria as asymptotically unbiased estimators of the natural risk function corresponding to the loss function used to obtain SDID estimators. By using these criteria, the dimension of $\bm{\theta}$ is determined and conponent of $\bm{x}$ is selected. As its natural risk function, based on (4), we define a weighted risk function,
@@ 4.1 Risk function and its asymptotic evaluation in the semiparametric difference-in-difference method
According to the derivation of the conventional $C_{p}$ criterion, we remove the expectation in the first term, ignore the second term which is independent of the models, and asymptotically evaluate the third term. For simplicity, we refer to the third term as the bias when the risk is simply estimated. If we define $\varepsilon_{i}\equiv\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha})\Delta_{i}-a(\bm{x}_{i})$ using $a(\bm{x}_{i})$ in (1), then the third term can be written as
@@
$=2\sum_{i=1}^{n}{\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\{\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}-a(\bm{x}_{i})\}]$
$=2\sum_{i=1}^{n}{\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\bm{x}_{i}^{{\rm T}}(\hat{\bm{\theta}}-\bm{\theta}^{*})].$ (17)
The second equality follows from ${\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})a(\bm{x}_{i})]=0$ and ${\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\allowbreak\bm{x}^{{\rm T}}_{i}\bm{\theta}^{*}]=0$ since ${\rm E}[\varepsilon_{i}\mid\bm{x}_{i}]=0$ holds from (3). Here, we use the same $e^{[1]}(\bm{x}_{i};\bm{\alpha})$ as that used to construct $\hat{\bm{\theta}}$. Then, letting $c^{\rm limit}$ be the weak limit of the third term with expectation removed, we consider ${\rm E}[c^{\rm limit}]$ as the asymptotic evaluation of the third term of (16). Note that we consider a risk function based on the loss function in deriving the SDID estimator, but it is also possible to consider a simple squared error.
@@ 4.2 Derivation of Model Selection Criterion
In the case of known propensity scores, we denote $e^{[1]}(\bm{x}_{i};\bm{\alpha}^{*})=e^{[1]}(\bm{x}_{i})$, $\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha}^{*})=\rho(\bm{t}_{i},\bm{x}_{i})$ for simplicity. The main term of the error of the SDID estimator is given by
@@
This expectation for the case in which $i\neq j$ is the product of the expectations for $i$ and $j$ from the independence among samples. The expected value for $i$ can be calculated as
@@
Therefore, we have only to consider the case $i=j$. Since $a(\bm{x}_{j})-\bm{x}_{j}^{{\rm T}}\bm{\theta}^{*}$ contains only $\bm{x}_{j}$ as a random variable, this term disappears from 18, and the following theorem is obtained.
@@
When the propensity score is known,
${\rm E}[c^{\rm limit}]=2{\rm E}[\varepsilon e^{[1]}(\bm{x})\bm{x}^{{\rm T}}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}e^{[1]}(\bm{x})\bm{x}\varepsilon]=2{\rm tr}\{{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[\sigma^{2}(\bm{x})e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}]\}$
is obtained, where
@@
While empirical evaluation of ${\rm E}[c^{\rm limit}]$ may seem easily possible, addressing the term ${\rm E}[{\rm E}[\rho(\bm{t},\bm{x})\Delta\mid\bm{x}]^{2}e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}]$ must be considered. The reason is that this cannot be empirically evaluated as $1/n\sum_{i=1}^{n}\rho(\bm{t},\bm{x})^{2}\allowbreak\Delta^{2}e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}$. In this case, we will substitute $\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}$ for ${\rm E}[\rho(\bm{t}_{i},\bm{x}_{i})\Delta\mid\bm{x}_{i}]$. As a result, when the propensity score is known, we propose
@@
$+2{\rm tr}\biggl{(}\biggl{\{}\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\biggr{\}}^{-1}\biggl{[}\sum_{i=1}^{n}\Bigl{\{}\rho(\bm{t}_{i},\bm{x}_{i})^{2}\Delta^{2}-(\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}})^{2}\Bigr{\}}e^{[1]}(\bm{x}_{i})^{2}\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\biggr{]}\biggr{)}$ (20)
as the model selection criterion for the SDID method.
4.3 Extension of the QICw
In order to provide a comparator for the proposed model selection criterion, we extend the QICw (Platt et al. 2013) for the setting of this article, which is the model selection criterion for propensity score weighted estimation on missing data. First, for the term of goodness of fit, we construct it using the same ideas as for the QICw. Specifically, the log-likelihood function with no missing data is differentiated, a weighted potential outcome variable is substituted for the one whose expected value is the same as the outcome variable, the differentiation is reverted, and the estimator is substituted for the parameter. The number of parameters is used as the penalty term. In this setting, the log-likelihood function with no missing data should be
$-\frac{1}{2\sigma^{2}}\sum_{i=1}^{n}[\{\mbox{$y^{[1]}_{i}(1)-y^{[0]}_{i}(1)$ when $d_{i}^{[1]}=1$}\}-\bm{x}_{i}^{{\rm T}}\bm{\theta}]^{2},$
and if we take $\rho_{i}\Delta_{i}$ as the weighted potential outcome variable, the outcome variable and the conditional expectation will be the same from (3). Then, the sum of the goodness of fit and penalty terms multiplied by the constant $2\sigma^{2}$, that is,
@@ 4.3 Extension of the QICw
we call QICw for SDID. In fact, if $\sigma^{2}(\bm{x})$ is constant in our proposed criterion, then it is equivalent to the QICw. The problem is to estimate the variance $\sigma^{2}$ of “$y^{[1]}_{i}(1)-y^{[0]}_{i}(1)$ when $d_{i}^{[1]}=1$”. This estimated value cannot be obtained from the data using only the present assumptions. Since $y^{[1]}_{i}(1)-y^{[0]}_{i}(1)=\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$ holds from $y^{[0]}(0)=y^{[1]}(0)$, and from Assumption 1, the expected value of “$\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$ in $d^{[1]}_{i}=1$” is the same as the expected value of “$\Delta_{i}^{[1]}$ in $d^{[0]}_{i}=0$ minus $\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$ in $d^{[1]}_{i}=1$”. Since the variance must be evaluated, we further assume that $\Delta_{i}^{[1]}$ and $\Delta_{i}^{[0]}$ are independent conditioned on $d^{[1]}=1$ and $d^{[0]}=1$, respectively, and calculate that
@@
$={\rm E}[\Delta^{2}\mid d^{[1]}=1]-{\rm E}[\Delta\mid d^{[1]}=1]^{2}+{\rm E}[\Delta^{2}\mid d^{[0]}=1]-{\rm E}[\Delta\mid d^{[0]}=1]^{2}$
From the above, we define $n_{1}\equiv\sum_{i=1}^{n}d^{[1]}_{i}$ and $n_{0}\equiv\sum_{i=1}^{n}d^{[0]}_{i}$, and
@@
Let us compare $\sigma^{2}(\bm{x})$ in the proposed criterion with $\sigma^{2}$ in the QICw. In (19), if all expected values do not actually depend on $\bm{x}$, then
$\sigma^{2}(\bm{x})=$ ${\rm E}\bigg{[}\frac{\Delta^{[1]2}}{e^{[1]}}\ \bigg{|}\ d^{[1]}=1\bigg{]}-{\rm E}[\Delta^{[1]}\mid d^{[1]}=1]^{2}+{\rm E}\bigg{[}\frac{\Delta^{[0]2}}{e^{[0]}}\ \bigg{|}\ d^{[0]}=1\bigg{]}-{\rm E}[\Delta^{[0]}\mid d^{[0]}=1]^{2}$
@@ 4.4 Numerical experiments
By simulations, we evaluate the performance of the proposed criterion (20) when the propensity score is known. As data generation settings, we consider the Case 1-1,
@@
$y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+1+\beta^{*}x_{1}+\epsilon^{[1]}.$
In both cases, $\beta^{*}$ represents the degree of contribution of the covariate $x_{1}$ to the causal effect.
In Table 2, for some $\beta^{*}$ and sample size $n$, we examine whether the second term of (20), the penalty term of the proposed criterion, accurately approximates the bias (17) that we originally wanted to evaluate. In all cases, our penalty terms are close to the true values of the bias, while the penalty terms in the QICw significantly underestimate the bias.
Table 2: Bias evaluation when the propensity score is known: True is the true value of the bias evaluated by the Monte Carlo method, and Proposal is the penalty term of the proposed method. The penalty term of the QICw is also given as a reference.
@@
Next, we examine the performance of model selection using the proposed criteria. Data is generated from
@@
$y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+\bm{\beta}^{[1]*}{}^{{\rm T}}\bm{x}+\epsilon^{[1]},$
and we consider the Case 2-1,
@@
Table 3 compares the proposed criterion and the QICw by selecting models using the forward selection method. As the main index to measure the goodness of the criteria, we use the empirical estimator of the risk function, specifically, the average of 3000 calculations of
@@
In the Case 2-2 setting, where there are relatively few unnecessary covariates among the candidates, that is, the negative effect of false positives is less likely to appear, the QICw slightly exceeds the proposed criterion in some cases for $\beta^{*}=0.5$ and $\beta^{*}=1.0$. Otherwise, the proposed criterion outperforms the QICw, especially in settings with many covariates unrelated to the true structure such as Case 2-1 and Case 2-3, a clear difference is observed. Since the QICw underestimates the penalty term, the false positive value must be large and too many extraneous variables must be selected in all settings.
Table 3: Comparison of the proposed criterion and the QICw through model selection when the propensity score is known: Risk is the empirical estimate of the risk of the selected model, TP is the average number of covariates that are true positives, and FP is the average number of covariates that are false positives.
@@
5 Model Selection Criteria for the CBD Method

### v3

@@ 3.2 Numerical experiments
Let us examine the robustness of the proposed estimation against model misspecification through simulation studies. The setting for data generation follows
@@
Throughout this paper, we will suppose that all random numbers are generated independently. The parameter $\beta^{*}$ represents the magnitude of the contribution from the covariate $x_{1}$ to the causal effect. The propensity score is estimated using only $x_{1}$; that is, the parameter $\alpha^{*}$ represents the magnitude of the contribution from $x_{2}$ to the true propensity scores, which is not included in the model for estimating the propensity scores. Note that when $\alpha^{*}\neq 0$, $x_{2}$ should also be included in the model, and we are doing a model misspecification because it was not included. Anyway, since (7) holds, the proposed estimation is theoretically guaranteed to be consistent. In this setting, we will compare the proposed method with a naive method estimating the propensity score by the MLE and check whether the proposed method is robust to misspecification of the propensity score model. Note that each of the numerical experiments described below was repeated 3,000 times.
The true values of the ATT, the mean estimates, and their empirical confidence intervals are listed in Table 1 for several $\alpha^{*}$, $\beta^{*}$ and sample sizes $n$. With increasing $\alpha^{*}$, which represents the magnitude of the misspecification of the propensity score model, the bias of the naive method becomes larger, and the true value is eventually no longer included in the confidence interval. Regarding the proposed method, the approximation accuracy is considerably high when the identity matrix is used as the weighting matrix in the generalized method of moments. On the other hand, when the optimal matrix is used as the weighting matrix, although the accuracy is often comparable to that of using the identity matrix, in some cases, it is similar to that of the naive method. Since there are settings in simulation studies where the optimal matrix is close to being degenerate, those cases in which the accuracy deteriorates to that of the naïve method may be attributed to the fact that the estimation of $\theta$ is not stable. Regardless of whether (7) holds, if the propensity score model is estimated correctly, both the proposed method and the naive method can estimate the ATT without bias.
Table 1: Comparison of estimates of the ATT when the propensity scores are given by the CBD method and the MLE: True is the true value of the ATT, CBD-id is the CBD method when the weighting matrix in the generalized method of moments is the identity matrix, and CBD-opt is the CBD method when it is the optimal matrix. Est indicates the average of the estimates obtained by each simulation, and Conf Int indicates its empirical 95% confidence interval.
@@
4 Model selection criterion for semiparametric difference-in-differences when propensity scores are known
4.1 Risk function and its asymptotic evaluation in semiparametric difference-in-differences
As mentioned in Section 1, there are no reasonable model selection criteria even for the basic setting treated in Abadie (2005). To remedy this situation, in this section, we will first derive a model selection criterion for when the propensity scores are known and then develop a model selection criterion for when the proposed CBD method is used. Specifically, by using the ideas of Baba et al. (2017) and Baba and Ninomiya (2025), we will derive the model selection criteria as asymptotically unbiased estimators of the natural risk corresponding to the loss function used to obtain SDID estimators. Using these criteria, the dimension of $\bm{\theta}$ is determined and the component of $\bm{x}$ is selected. A natural choice is to define a weighted risk based on (4) as
@@ 4.1 Risk function and its asymptotic evaluation in semiparametric difference-in-differences
Following the derivation of the conventional $C_{p}$ criterion, we will remove the expectation in the first term, ignore the second term which is independent of the models, and asymptotically evaluate the third term. For simplicity, we will refer to the third term as a bias when the risk is simply estimated. If we define $\varepsilon_{i}\equiv\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha})\Delta_{i}-a(\bm{x}_{i})$ using $a(\bm{x}_{i})$ in (1), the third term can be written as
@@
$=2\sum_{i=1}^{n}{\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\{\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}-a(\bm{x}_{i})\}]=2\sum_{i=1}^{n}{\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\bm{x}_{i}^{{\rm T}}(\hat{\bm{\theta}}-\bm{\theta}^{*})].$ (16)
The second equality is satisfied by ${\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})a(\bm{x}_{i})]=0$ and ${\rm E}[\varepsilon_{i}e^{[1]}(\bm{x}_{i};\bm{\alpha})\allowbreak\bm{x}^{{\rm T}}_{i}\bm{\theta}^{*}]=0$ since it follows from (3) that ${\rm E}[\varepsilon_{i}\mid\bm{x}_{i}]=0$. Here, we have used the same $e^{[1]}(\bm{x}_{i};\bm{\alpha})$ as that was used to construct $\hat{\bm{\theta}}$. Then, letting $c^{\rm limit}$ be the weak limit of the third term with the expectation removed, we consider ${\rm E}[c^{\rm limit}]$ as the asymptotic evaluation of the third term of (15). Note that while we consider a risk based on the loss function in deriving the SDID estimator, it is also possible to consider a simple squared error.
@@ 4.2 Derivation of model selection criterion
In the case of known propensity scores, we will denote $e^{[1]}(\bm{x}_{i};\bm{\alpha}^{*})$ as $e^{[1]}(\bm{x}_{i})$ and $\rho(\bm{t}_{i},\bm{x}_{i};\bm{\alpha}^{*})$ as $\rho(\bm{t}_{i},\bm{x}_{i})$ for simplicity. The main term of the error of the SDID estimator can be expressed as
@@
This expectation for the case of $i\neq j$ is the product of the expectations for $i$ and $j$, from the independence among the samples. The expectation for $i$ can be calculated as
@@
Therefore, we only have to consider the case of $i=j$. Since $a(\bm{x}_{j})-\bm{x}_{j}^{{\rm T}}\bm{\theta}^{*}$ contains only $\bm{x}_{j}$ as a random variable, this term disappears from (17), and the following theorem is obtained.
@@
When the propensity scores are known, it holds that
${\rm E}[c^{\rm limit}]=2{\rm E}[\varepsilon e^{[1]}(\bm{x})\bm{x}^{{\rm T}}{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}e^{[1]}(\bm{x})\bm{x}\varepsilon]$
$=2{\rm tr}\{{\rm E}[e^{[1]}(\bm{x})\bm{x}\bm{x}^{{\rm T}}]^{-1}{\rm E}[\sigma^{2}(\bm{x})e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}]\},$
where
@@
Although one may think that an empirical evaluation of ${\rm E}[c^{\rm limit}]$ is easy, we must consider how to deal with the term ${\rm E}[{\rm E}[\rho(\bm{t},\bm{x})\Delta\mid\bm{x}]^{2}e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}]$. This cannot be empirically evaluated as $n^{-1}\sum_{i=1}^{n}\rho(\bm{t},\bm{x})^{2}\allowbreak\Delta^{2}e^{[1]}(\bm{x})^{2}\bm{x}\bm{x}^{{\rm T}}$. Here, we will substitute $\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}}$ into ${\rm E}[\rho(\bm{t}_{i},\bm{x}_{i})\Delta\mid\bm{x}_{i}]$. As a result, when the propensity scores are known, we propose
@@
$+2{\rm tr}\biggl(\biggl\{\sum_{i=1}^{n}e^{[1]}(\bm{x}_{i})\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\biggr\}^{-1}\biggl[\sum_{i=1}^{n}\{\rho(\bm{t}_{i},\bm{x}_{i})^{2}\Delta^{2}-(\bm{x}_{i}^{{\rm T}}\hat{\bm{\theta}})^{2}\}e^{[1]}(\bm{x}_{i})^{2}\bm{x}_{i}\bm{x}_{i}^{{\rm T}}\biggr]\biggr)$ (19)
as the model selection criterion for the basic SDID approach.
4.3 Extension of QICw
As a comparator for the proposed model selection criterion, we will extend QICw (Platt et al. 2013), which is an intuitive model selection criterion for weighted estimation using propensity scores, to the setting of this paper. First, we construct the goodness-of-fit term using the same idea as for QICw. To do so, the log-likelihood function for full data is differentiated, a weighted potential outcome variable is substituted into the one whose expectation is the same as the outcome variable, the differentiation is reverted, and the estimator is substituted into the parameter. The number of parameters is used as the penalty term. Within this setting, the log-likelihood function for full data is
$-\frac{1}{2\sigma^{2}}\sum_{i=1}^{n}[\{\mbox{$y^{[1]}_{i}(1)-y^{[0]}_{i}(1)$ when $d_{i}^{[1]}=1$}\}-\bm{x}_{i}^{{\rm T}}\bm{\theta}]^{2}.$
If we take $\rho_{i}\Delta_{i}$ as the weighted potential outcome variable, its conditional expectation can be shown to be the same as the conditional expectation of the outcome variable from (3). We will call the sum of the goodness-of-fit term and the penalty term multiplied by a constant $2\sigma^{2}$, that is,
@@ 4.3 Extension of QICw
the QICw for the SDID approach. In fact, if $\sigma^{2}(\bm{x})$ is a constant, then our proposed criterion is equivalent to the QICw. The problem is to estimate the variance $\sigma^{2}$ of the difference, $y^{[1]}_{i}(1)-y^{[0]}_{i}(1)$ when $d_{i}^{[1]}=1$. This estimate cannot be obtained from the data using only the present assumptions. Since it follows from $y^{[0]}(0)=y^{[1]}(0)$ that $y^{[1]}_{i}(1)-y^{[0]}_{i}(1)=\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$, and from Assumption 1, the expectation of $\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$ in $d^{[1]}_{i}=1$ is the same as the expectation of $\Delta_{i}^{[1]}$ in $d^{[0]}_{i}=0$ minus $\Delta_{i}^{[1]}-\Delta_{i}^{[0]}$ in $d^{[1]}_{i}=1$. However, in order to evaluate the variance, we will assume that $\Delta_{i}^{[1]}$ and $\Delta_{i}^{[0]}$ are independent conditioned on $d^{[1]}=1$ and $d^{[0]}=1$, from which we have
@@
$={\rm E}[\Delta^{2}\mid d^{[1]}=1]-{\rm E}[\Delta\mid d^{[1]}=1]^{2}+{\rm E}[\Delta^{2}\mid d^{[0]}=1]-{\rm E}[\Delta\mid d^{[0]}=1]^{2}.$
Moreover, defining $n_{1}\equiv\sum_{i=1}^{n}d^{[1]}_{i}$ and $n_{0}\equiv\sum_{i=1}^{n}d^{[0]}_{i}$, we will use
@@
Let us compare $\sigma^{2}(\bm{x})$ in the proposed criterion with $\sigma^{2}$ in the QICw. In (18), if all expectations do not actually depend on $\bm{x}$, then
$\sigma^{2}(\bm{x})=\$${\rm E}\bigg[\frac{\Delta^{[1]2}}{e^{[1]}}\ \bigg|\ d^{[1]}=1\bigg]-{\rm E}[\Delta^{[1]}\mid d^{[1]}=1]^{2}+{\rm E}\bigg[\frac{\Delta^{[0]2}}{e^{[0]}}\ \bigg|\ d^{[0]}=1\bigg]-{\rm E}[\Delta^{[0]}\mid d^{[0]}=1]^{2}$
@@ 4.4 Numerical experiments
Here, the performance of the proposed criterion (19) when the propensity scores are known is evaluated in simulation studies. The data generation settings include Case 1-1,
@@
$y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+1+\beta^{*}(x_{1}+x_{2})+\epsilon^{[1]}.$
In both cases, $\beta^{*}$ represents the magnitude of contribution of the covariate $x_{1}$ to the causal effect.
Table 2 shows whether the second term in (19), the penalty term of the proposed criterion, accurately approximates the bias in (16) that we originally wanted to evaluate. For all values of $\beta^{*}$ and sample size $n$, our penalty terms are close to the true values of the bias, while the evaluated penalty terms in the QICw significantly underestimate the bias.
Table 2: Bias evaluation for the case where the propensity scores are known: True denotes the true value of the bias evaluated by the Monte Carlo method, and Proposal the evaluated penalty term of the proposed criterion. The evaluated penalty term of the QICw is listed as a reference.
@@
Next, let us examine the performance of the model selection using the proposed criterion, with data generated from
@@
$y^{[0]}(1)=y^{[0]}(0)+\epsilon^{[0]},\quad y^{[1]}(1)=y^{[1]}(0)+\bm{x}^{{\rm T}}\bm{\beta}^{[1]*}+\epsilon^{[1]},$
in Case 2-1,
@@
Table 3 compares the proposed criterion and the QICw when using the forward selection method. As the main index to measure the goodness of the criteria, we use an empirical estimate of the risk, specifically, the average of 3,000 values of
@@
In Case 2-2, where there are relatively few unnecessary covariates among the candidates, that is, where the negative effect of false positives is less likely to appear, the QICw slightly outperforms the proposed criterion for $\beta^{*}=0.5$ and $\beta^{*}=1.0$. However, the proposed criterion outperforms the QICw for the other values of beta in Case 2-2, and it outperforms the QICw for all values in Case 2-1 and Case 2-3, i.e., settings with many covariates are unrelated to the true structure. Here, since the QICw underestimates the penalty term, the false positive must be large and too many covariates are selected in all settings.
Table 3: Comparison of the proposed criterion and the QICw through model selection for case in which the propensity scores are known: Risk denotes the empirical estimate of the risk of the selected model, TP the average number of covariates that are true positive, and FP the average number of covariates that are false positive.
@@
5 Model selection criterion for covariate balancing difference-in-differences